# MEMOIR: Lifelong Model Editing with Minimal Overwrite and Informed Retention for LLMs

### v1

@@ Abstract
Language models deployed in real-world systems often require post-hoc updates to incorporate new or corrected knowledge. However, editing such models efficiently and reliably—without retraining or forgetting previous information—remains a major challenge. Existing methods for lifelong model editing either compromise generalization, interfere with past edits, or fail to scale to long editing sequences. We propose MEMOIR, a novel scalable framework that injects knowledge through a residual memory, i.e., a dedicated parameter module, while preserving the core capabilities of the pre-trained model. By sparsifying input activations through sample-dependent masks, MEMOIR confines each edit to a distinct subset of the memory parameters, minimizing interference among edits. At inference, it identifies relevant edits by comparing the sparse activation patterns of new queries to those stored during editing. This enables generalization to rephrased queries by activating only the relevant knowledge while suppressing unnecessary memory activation for unrelated prompts. Experiments on question answering, hallucination correction, and out-of-distribution generalization benchmarks across LLaMA-3 and Mistral demonstrate that MEMOIR achieves state-of-the-art performance across reliability, generalization, and locality metrics, scaling to thousands of sequential edits with minimal forgetting.
@@ 1 Introduction
Large language models (LLMs) exhibit strong performance across a wide range of tasks, enabled by large-scale pre-training on diverse and extensive datasets (gpt-3, ; Chowdhery2022PaLMSL, ). However, LLMs can generate outdated or inaccurate information and can perpetuate biases during deployment llm_factual_error ; hallucination_survey ; chatgpt_bias ; llm_factual_error , which requires further continuous updates to their knowledge base after pre-training to correct these behaviors. Although fine-tuning on new corpora can revise model behavior, it is computationally expensive and prone to catastrophic forgetting  (Goodfellow2013AnEI, ; Oswald2019ContinualLW, ). These limitations motivate an emerging line of work known as lifelong model editing (grace, ; wang2024wise, ), which aims to continuously update the model’s knowledge in an efficient and localized manner.
@@
Figure 1: Left: trade-offs among reliability (Rel.), generalization (Gen.), and locality (Loc.) for 1000 continual edits. Right: Average performance of Rel., Gen., and Loc. under varying numbers of edits. Both evaluated on LLaMA-3-8B-Instruct (llama-3, ) with ZsRE dataset (zsre, ). MEMOIR delivers best balance among the three metrics and scale with large number of edits.
In this paper, we propose MEMOIR, (Model Editing with Minimal Overwrite and Informed Retention), which delivers a desirable balance between reliability, generalization, and locality for large number of edits (Figure 1). MEMOIR introduces a memory module, i.e., a dedicated fully-connected layer in a single transformer block where all edits are performed. MEMOIR mitigates catastrophic forgetting by allocating distinct parameter subsets to each edit and retrieving them during inference to activate only the relevant knowledge for a given prompt. During editing, we propose a structured sparsification of the input with sample-dependent masks to this layer, dynamically activating only a sample-specific subset of parameters in the introduced memory module in the forward pass. MEMOIR therefore promotes the distribution of new knowledge across the parameter space, reducing the overwriting of previous updates and significantly mitigating catastrophic forgetting. During inference, we use the sparsification pattern of a prompt to infer if it semantically corresponds to an edited prompt, and route the activations accordingly. This targeted knowledge activation eliminates the need for large corpora of irrelevant samples to regularize training (memit, ; fang2025alphaedit, ; wang2024wise, ). Specifically, if the input corresponds to a rephrased version of an edit, we activate only the relevant knowledge of that edit to align the representation of the prompt. Conversely, the introduced module is deactivated when detecting irrelevant prompts, preserving the pre-trained knowledge of the LLM.
@@
  • We extensively evaluate MEMOIR on Q&A, hallucination correction, and OOD generalization on temporal tasks, demonstrating state-of-the-art results across LLaMA-3 (llama-3, ), Mistral (mistral, ), LLaMA-2 (llama-2, ) and GPT-J (gpt-j, ) architectures compared to all previous methods.
  • We extend the previous state-of-the-art edit horizon to 7000 edits and show that MEMOIR consistently delivers superior performance in the challenging setting of sequential singular edits compared with previous editing methods.
@@ 2 Lifelong Model Editing
LLMs are based on the transformer architecture vaswani2017attention and consist of $L$ identical blocks containing multi-head attention modules followed by feed-forward network (FFN) modules. For each block $\ell\in[L]$, the FFN module consists of two fully connected layers, parameterized by $\bm{W}_{\text{fc}}^{\ell}$ and $\bm{W}_{\text{proj}}^{\ell}$, and an activation function. Previous work shows that knowledge is typically stored in the middle blocks of an LLM rome , and updating the parameters of a single middle block rather than fine-tuning the entire model can be effective to edit factual knowledge wang2024wise ; grace . In what follows, to lighten notation, we will omit the layer-indexing superscript. Let $\bm{h}$ represent the output from the attention module; then the output of the FFN module is formally defined as:
@@
for an element-wise activation function $\sigma(\cdot)$. The FFN module can be interpreted as a two-layer key-value memory system (geva-etal-2021-transformer, ), where the second layer $\bm{W}_{\text{proj}}$ functions as a linear associative memory (rome, ), mapping a sequence of input vectors $[\bm{a}_{1},\bm{a}_{2},\dots]$ to corresponding output vectors $[\bm{v}_{1},\bm{v}_{2},\dots]$. This memory-like behavior enables information retrieval, where $\bm{a}=\sigma(\bm{W}_{\text{fc}}\ \bm{h})$ and $\bm{v}=\text{FFN}(\bm{h})$ respectively represent the input and output vectors of the layer $\bm{W}_{\text{proj}}$. Leveraging this perspective, previous works have shown the effectiveness of modifying the output of $\bm{W}_{\text{proj}}$ to update the model’s stored knowledge (rome, ; geva-etal-2021-transformer, ; wang2024wise, ). However, continuously updating $\bm{W}_{\text{proj}}$ in lifelong editing gradually overwrites previous edits, leading to significant forgetting and severe performance degradation with a large number of edits, as shown in Figure 1.
Figure 2: (a) Overall framework of MEMOIR during inference stage. The edited output combines the outputs of the original layer output and the residual memory. The input to the residual memory conditionally activates specific columns in the residual memory to retrieve relevant knowledge. (b) During editing, each edit modifies only a designated subset of columns in the residual memory, minimizing overwriting to previous edits in the memory. For visualization we transpose the weight matrices.
@@ 3 Memory Editing with Minimal Overwrite and Informed Retention
We propose MEMOIR: Model Editing with Minimal Overwrite and Informed Retention. During editing, MEMOIR distributes knowledge storage by allocating distinct parameter subsets for each edit to reduce overwriting of previous edits (Section 3.1). At inference stage, MEMOIR dynamically activates relevant parameters related to the knowledge of each input prompt (Section 3.2). Through this design, MEMOIR achieves a superior balance among reliability, generalization and locality even with a large number of edits. We start by presenting the general framework of MEMOIR illustrated in Figure 2.
@@ 3.1 Editing: distributing knowledge across memory
Prior parametric approaches on knowledge editing memit ; fang2025alphaedit typically store the knowledge of each edit across the entire trainable parameter space. However, unconstrained update on all trainable parameters lead to severe catastrophic forgetting, as new edits quickly overwrite knowledge from previous ones, as shown in Figure 1. In this work, we draw motivation from the continual learning literature leveraging sparsity mallya2018packnet ; wortsman2020supermasks ; mirzadeh2020understanding ; saha2021gradient ; wang2022sparcl ; abbasi2022sparsity to mitigate catastrophic forgetting, and propose a structured sparsification of input activations to the residual memory during the forward pass. This ensures that each edit modifies only a subset of the trainable parameters, promoting less overlap in representations and leading to more localized updates across different edits.
@@
Applying the sparse mask $\mathcal{M}(\bm{a}(\bm{x}))$ to the input activations retains only $k\ll D$ active indices, setting the rest to zero. As a result, gradient updates are restricted to the $k$ corresponding columns in $\bm{W}_{\text{m}}$, where the knowledge for input $\bm{x}$ is explicitly stored. The other columns remain unchanged, effectively preserving previously stored edits and minimizing interference. Since the last token captures the most informative summary of the input, the mask is derived from its activation and applied across all tokens.
@@
A crucial aspect of this sparse masking is the selection mechanism for the active indices, which intuitively should satisfy two criteria: selection of diverse features and semantic coherence. Diversity in feature selection encourages the model to consider a broader set of features, reducing the risk of overfitting on dominant features and mitigating catastrophic forgetting by distributing updates across the trainable parameters of the memory module. Semantic coherence ensures that semantically similar input prompts yield similar selected masks, facilitating the activation of relevant knowledge in the memory module for previously unseen yet semantically related queries. Therefore, we opt for structured sparsity meier2008group .
The masking should be sample-dependent so that different feature subsets are chosen based on the input characteristics. A naive way to introduce diversity in feature selection is to generate a random seed for index selection directly from the input activations; however, this results in highly input-specific masks that fail to generalize to semantically similar prompts, such as rephrased edits, defeating the purpose of structured and reusable knowledge storage. To address this limitation, we turn to a more robust fingerprinting mechanism based on the inherent structure of the activations themselves.
@@
For an input $\bm{a}=\bm{a}(\bm{x})\in\mathbb{R}^{D}$, we compute a $\mathrm{top-k}$ selection based on feature magnitudes, mapping $\bm{a}$ to a binary mask $\mathcal{T}(\bm{a})$. Denoting as $a_{(k)}$ the $k$-th largest value using order statistics notation, the mask is formally defined as $\mathcal{T}(\bm{a})_{j}=\mathds{1}[a_{j}\geq a_{(k)}],\;j\in[D]$. This strategy ensures that semantically similar input prompts—naturally producing similar activation patterns—yield similar masks, directly addressing the generalization shortcoming of random seeds. Given the structured and highly correlated nature of activations in LLMs, this step effectively acts as a Locality-Sensitive Hashing (LSH) indyk1998approximate mechanism, mapping semantically similar input prompts to identical or closely related binary codes norouzi2012hamming .
Selecting the exact top-$k$ features concentrates updates on the most salient and frequently shared features across input prompts. Due to their large activation magnitudes, they are prune to be sensitive to weight changes and updates to the corresponding parameters increase the likelihood of interference with previously acquired knowledge and exacerbate catastrophic forgetting. To introduce diversity without sacrificing semantic coherence, we introduce a form of controlled randomness by applying a fixed random permutation $\pi:[D]\mapsto[D]$ to the mask, decoupling the selection of salient features from their fixed positions. This strategically redirects updates away from the most influential features and onto less critical ones. Semantic coherence is maintained—since similar prompts still yield overlapping top-k indices under the fixed permutation—while diversifying the actual features activated in the residual memory. Due to the information redundancy in the LLM activations and the large number of active indices $k$, the application of the random permutation $\pi$ strikes a good balance between diversity and information retention. We ablate its importance in Table 11 of the appendix, showing that both the top-k selection and further permutation are crucial to the performance.
@@ 3.2 Inference: activating knowledge conditional to inference prompt
After editing, the model stores the updated knowledge in the memory module $\bm{W}_{\text{m}}$, while the pre-trained knowledge is preserved in the original layer $\bm{W}_{0}$. To correctly respond to diverse prompts at inference stage, the model needs to retrieve the relevant knowledge from its memory. To do so, we propose the following two-step process (i) identifying the type of
@@
prompt during inference—whether it is an edited prompt, a rephrased version of an edited prompt, or a prompt irrelevant to the edits; and (ii) activating only the relevant knowledge in the memory. Specifically, it should retrieve the relevant knowledge distributed in corresponding combination of columns in $\bm{W}_{\text{m}}$ for the edited prompts and their rephrased variants, while deactivating the entire residual memory for irrelevant prompts for preserving locality.
During editing, we construct a database of sparse masks of each edited sample. Let $\bm{x}\in\mathcal{D}_{\textrm{edit}}$ be an edited prompt. Given that the layers prior to the introduced memory module are frozen, the application of the TopHash mechanism will yield the same sparse mask $\mathcal{M}(\bm{a}(\bm{x}))$ during both editing and inference, allowing the reliable identification of the prompt as part of the edit set by finding the exact match in the above database. Since transformers cluster semantically similar input prompts together due to its ability to generalize across paraphrases sentencebert , we expect that $\bm{x}$ should have activations, as well as TopHash masks, with high similarity w.r.t. its semantically rephrased variants. Conversely, semantically unrelated prompts will exhibit low similarity in both activation and mask space to the edited prompt $\bm{x}$.
@@
$\bm{x}_{\mathrm{match}}$ $=\operatorname*{arg\,min}_{\bm{x}^{\prime}\in\mathcal{D}_{\mathrm{edit}}}d_{\mathrm{H}}\bigl{(}\mathcal{M}(\bm{a}(\bm{x})),\,\mathcal{M}(\bm{a}(\bm{x}^{\prime}))\bigr{)}$
We define the optimal overlapping ratio [^1] Note that the overlapping ratio is equivalent to the Hamming distance for two binary vectors with the exact same number of zeroes and ones. between a prompt $\bm{x}$ and its closest edit $\bm{x}_{\textrm{match}}$, as:
@@
We intuitively classify the prompt type of the input $\bm{x}$ based on $R_{\mathrm{match}}$ and a threshold $\tau$ and activate only the relevant knowledge from memory accordingly. Formally, the edited feedforward output at inference is then computed as:
$\text{FFN}_{\text{edited}}(\bm{a}(\bm{x}))=\begin{cases}\bm{W}_{0}\ \bm{a}(\bm{x})+\bm{W}_{\text{m}}\left(\mathcal{M}(\bm{a}(\bm{x}_{\mathrm{match}}))\odot\bm{a}(\bm{x})\right)&\text{if }R_{\mathrm{match}}(\bm{x})\in[\tau,1.0],\\ \bm{W}_{0}\ \bm{a}(\bm{x})&\text{if }R_{\mathrm{match}}(\bm{x})\in[0,\tau)\end{cases}$ (4)
Specifically, if $R_{\mathrm{match}}(\bm{x})\geq\tau$, we identify $\bm{x}$ as semantically relevant to an edited prompt $\bm{x}_{\mathrm{match}}$, and apply the mask $\mathcal{M}(\bm{a}(\bm{x}_{\mathrm{match}}))$ to its activations to activate the columns in $\bm{W}_{\text{m}}$ storing the relevant knowledge. This mechanism ensures that semantically rephrased prompts are aligned with their original edited counterparts in representation space, enabling precise retrieval of the relevant residual memory. As a result, it significantly improves generalization to paraphrased queries. Meanwhile, the residual memory is deactivated for prompts considered irrelevant to all edited prompts, i.e., $R_{\mathrm{match}}(\bm{x})<\tau$, effectively preserving locality by relying solely on the pre-trained knowledge.
Prior approaches often require access to large corpora of irrelevant samples to enhance locality; either by sampling a different irrelevant sample during each edit (wang2024wise, ) or by performing a forward pass on a corpora of up to 100,000 samples to regularize editing (memit, ; fang2025alphaedit, ). MEMOIR, on the other hand, eliminates the need for such irrelevant samples during editing and offers a more flexible alternative, as it relies solely on the introduced conditional knowledge activation mechanism in Equation 4 to decide whether to use the edited or pre-trained knowledge.

### v2

@@ Abstract
Language models deployed in real-world systems often require post-hoc updates to incorporate new or corrected knowledge. However, editing such models efficiently and reliably—without retraining or forgetting previous information—remains a major challenge. Existing methods for lifelong model editing either compromise generalization, interfere with past edits, or fail to scale to long editing sequences. We propose MEMOIR, a novel scalable framework that injects knowledge through a residual memory, i.e., a dedicated parameter module, while preserving the core capabilities of the pre-trained model. By sparsifying input activations through sample-dependent masks, MEMOIR confines each edit to a distinct subset of the memory parameters, minimizing interference among edits. At inference, it identifies relevant edits by comparing the sparse activation patterns of new queries to those stored during editing. This enables generalization to rephrased queries by activating only the relevant knowledge while suppressing unnecessary memory activation for unrelated prompts. Experiments on question answering, hallucination correction, and out-of-distribution generalization benchmarks for LLaMA-3 and Mistral backbones demonstrate that MEMOIR achieves state-of-the-art performance across reliability, generalization, and locality metrics, scaling to thousands of sequential edits with minimal forgetting.
@@ 1 Introduction
Large language models (LLMs) exhibit strong performance across a wide range of tasks, enabled by large-scale pre-training on diverse and extensive datasets (gpt-3, ; Chowdhery2022PaLMSL, ). However, LLMs can generate outdated or inaccurate information and can perpetuate biases during deployment llm_factual_error ; hallucination_survey ; chatgpt_bias , which requires further continuous updates to their knowledge base after pre-training to correct these behaviors. Although fine-tuning on new corpora can revise model behavior, it is computationally expensive and prone to catastrophic forgetting  (Goodfellow2013AnEI, ; Oswald2019ContinualLW, ). These limitations motivate an emerging line of work known as lifelong model editing (grace, ; wang2024wise, ), which aims to continuously update the model’s knowledge in an efficient and localized manner.
@@
Figure 1: Left: trade-offs among reliability (Rel.), generalization (Gen.), and locality (Loc.) for 1,0001{,}000 continual edits. Right: Average performance of Rel., Gen., and Loc. under varying numbers of edits. Both evaluated on LLaMA-3-8B-Instruct (llama-3, ) with ZsRE dataset (zsre, ). MEMOIR delivers the best balance among the three metrics and scales with a large number of edits.
In this paper, we propose MEMOIR, (Model Editing with Minimal Overwrite and Informed Retention), which delivers a desirable balance between reliability, generalization, and locality for a large number of edits (Figure 1). MEMOIR introduces a memory module, i.e., a dedicated fully-connected layer in a single transformer block where all edits are performed. MEMOIR mitigates catastrophic forgetting by allocating distinct parameter subsets to each edit and retrieving them during inference to activate only the relevant knowledge for a given prompt. During editing, we propose a structured sparsification of the input with sample-dependent masks to this layer, dynamically activating only a sample-specific subset of parameters in the introduced memory module in the forward pass. MEMOIR therefore promotes the distribution of new knowledge across the parameter space, reducing the overwriting of previous updates and significantly mitigating catastrophic forgetting. During inference, we use the sparsification pattern of a prompt to infer if it semantically corresponds to an edited prompt, and route the activations accordingly. This targeted knowledge activation eliminates the need for large corpora of irrelevant samples to regularize training (memit, ; fang2025alphaedit, ; wang2024wise, ). Specifically, if the input corresponds to a rephrased version of an edit, we activate only the relevant knowledge of that edit to align the representation of the prompt. Conversely, the introduced module is deactivated when detecting irrelevant prompts, effectively preserving the pre-trained knowledge of the LLM.
@@
  • We extensively evaluate MEMOIR on Q&A, hallucination correction, OOD generalization, and multi-hop reasoning tasks, demonstrating state-of-the-art results across LLaMA-3 (llama-3, ), Mistral (mistral, ), LLaMA-2 (llama-2, ) and GPT-J (gpt-j, ) architectures compared to all previous methods.
  • We extend the previous edit horizon to 15,00015{,}000 edits and show that MEMOIR consistently delivers superior performance in the challenging setting of sequential singular edits compared with previous editing methods.
@@ 2 Lifelong Model Editing
LLMs are typically based on the transformer architecture vaswani2017attention and consist of $L$ identical blocks containing multi-head attention modules followed by feed-forward network (FFN) modules. For each block $\ell\in[L]$, the FFN module consists of two fully connected layers, parameterized by $\bm{W}_{\text{fc}}^{\ell}$ and $\bm{W}_{\text{proj}}^{\ell}$, and an activation function. Previous work shows that knowledge is typically stored in the middle blocks of an LLM rome , and updating the parameters of a single middle block rather than fine-tuning the entire model can be effective to edit factual knowledge wang2024wise ; grace . In what follows, to lighten notation, we will omit the layer-indexing superscript. Let $\bm{h}$ represent the output from the attention module; then the output of the FFN module is formally defined as:
@@
for an element-wise activation function $\sigma(\cdot)$. The FFN module can be interpreted as a two-layer key-value memory system (geva-etal-2021-transformer, ), where the second layer $\bm{W}_{\text{proj}}$ functions as a linear associative memory (rome, ), mapping a sequence of input vectors $\bm{a}_{1},\bm{a}_{2},\dots]$ to corresponding output vectors $\bm{v}_{1},\bm{v}_{2},\dots]$. This memory-like behavior enables information retrieval, where $\bm{a}=\sigma(\bm{W}_{\text{fc}}\ \bm{h})$ and $\bm{v}=\text{FFN}(\bm{h})$ respectively represent the input and output vectors of the layer $\bm{W}_{\text{proj}}$. Leveraging this perspective, previous works have shown the effectiveness of modifying the output of $\bm{W}_{\text{proj}}$ to update the model’s stored knowledge (rome, ; geva-etal-2021-transformer, ; wang2024wise, ; memit, ; fang2025alphaedit, ). However, continuously updating $\bm{W}_{\text{proj}}$ in lifelong editing gradually overwrites previous edits, leading to significant forgetting and severe performance degradation with a large number of edits, as shown in Figure 1.
Figure 2: (a) Overall framework of MEMOIR during inference stage. The edited output combines the outputs of the original layer output and the residual memory. The input to the residual memory conditionally activates specific columns in the residual memory to retrieve relevant knowledge. (b) During editing, each edit modifies only a designated subset of columns in the residual memory, minimizing overwriting of previous edits in the memory. For visualization, we transpose the weight matrices.
@@ 3 Memory Editing with Minimal Overwrite and Informed Retention
We propose MEMOIR: Model Editing with Minimal Overwrite and Informed Retention. During editing, MEMOIR distributes knowledge storage by allocating distinct parameter subsets for each edit to reduce overwriting of previous edits (Section 3.1). At inference stage, MEMOIR dynamically activates relevant parameters related to the knowledge of each input prompt (Section 3.2). Through this design, MEMOIR achieves a superior balance among reliability, generalization, and locality even with a large number of edits. We start by presenting the general framework of MEMOIR illustrated in Figure 2.
@@ 3.1 Editing: distributing knowledge across memory
Prior parametric approaches on knowledge editing memit ; fang2025alphaedit typically store the knowledge of each edit across the entire trainable parameter space. However, unconstrained updates on all trainable parameters lead to severe catastrophic forgetting, as new edits quickly overwrite knowledge from previous ones, as shown in Figure 1. In this work, we draw motivation from the continual learning literature leveraging sparsity mallya2018packnet ; wortsman2020supermasks ; mirzadeh2020understanding ; saha2021gradient ; wang2022sparcl ; abbasi2022sparsity to mitigate catastrophic forgetting, and propose a structured sparsification of input activations to the residual memory during the forward pass. This ensures that each edit modifies only a subset of the trainable parameters, promoting less overlap in representations and leading to more localized updates across different edits.
@@
Applying the sparse mask $\mathcal{M}(\bm{a}(\bm{x}))$ to the input activations retains only $k\ll D$ active indices, setting the rest to zero. As a result, gradient updates are restricted to the $k$ corresponding columns in $\bm{W}_{\text{m}}$, where the knowledge for input $\bm{x}$ is explicitly stored. The other columns remain unchanged, effectively preserving previously stored edits and minimizing interference. In practice, we construct the mask based on the activation averaged across all tokens in the prompt $\bm{x}$ and then apply it uniformly to the activation of each token. To suppress features with consistently high magnitude across all prompts, the average activations are centered by removing their mean computed with 100 irrelevant prompts. This enables a more diverse selection of masks across prompts. We provide more details and ablations in Appendix B.2. Next, we describe how the mask is derived using the TopHash mechanism.
@@
A fundamental aspect of this sparse masking is the selection mechanism for the active indices. Intuitively, this mechanism should satisfy two different criteria: selection of feature diversity and semantic coherence. Diversity in feature selection encourages the model to consider a broader set of features, reducing the risk of overfitting on dominant features and mitigating catastrophic forgetting by distributing updates across the trainable parameters of the memory module. Semantic coherence ensures that semantically similar input prompts yield similar selected masks, facilitating the activation of relevant knowledge in the memory module for previously unseen yet semantically related queries. Therefore, we opt for structured sparsity meier2008group .
The masking should be sample-dependent so that different feature subsets are chosen based on the input characteristics. A naive way to introduce diversity in feature selection is to generate a random seed for index selection directly from the input activations; however, this results in highly input-specific masks that fail to generalize to semantically similar prompts, such as rephrased edits,
@@
defeating the purpose of structured and reusable knowledge storage. To address this limitation, we turn to a more robust fingerprinting mechanism based on the inherent structure of the activations themselves. For an input $\bm{a}=\bm{a}(\bm{x})\in\mathbb{R}^{D}$, we compute a top-k selection based on feature magnitudes, mapping $\bm{a}$ to a binary mask $\mathcal{T}(\bm{a})$. Denoting as $a_{(k)}$ the $k$-th largest value using order statistics notation, the mask is formally defined as $\mathcal{T}(\bm{a})_{j}=\mathds{1}[a_{j}\geq a_{(k)}],\;j\in[D]$. This strategy ensures that semantically similar input prompts—naturally producing similar activation patterns—yield similar masks, directly addressing the generalization shortcoming of random seeds. Given the structured and highly correlated nature of activations in LLMs, this step effectively acts as a Locality-Sensitive Hashing (LSH) indyk1998approximate mechanism, mapping semantically similar input prompts to identical or closely related binary codes norouzi2012hamming .
Selecting the exact top-$k$ features concentrates updates on the most salient and frequently shared features across input prompts. Due to their large activation magnitudes, they are prone to being sensitive to weight changes and updates to the corresponding parameters increase the likelihood of interference with previously acquired knowledge and exacerbate catastrophic forgetting. To introduce diversity without sacrificing semantic coherence, we introduce a form of controlled randomness by applying a fixed random permutation $\pi:[D]\mapsto[D]$ to the mask, decoupling the selection of salient features from their fixed positions. This strategically redirects updates away from the most influential features and onto less critical ones. Semantic coherence is maintained—since similar prompts still yield overlapping top-k indices under the fixed permutation—while diversifying the actual features activated in the residual memory. Due to the information redundancy in the LLM activations and the large number of active indices $k$, the application of the random permutation $\pi$ strikes a good balance between diversity and information retention. We ablate its importance in Table 12 of the appendix, showing that both the top-k selection and further permutation are crucial to the performance.
@@ 3.2 Inference: activating knowledge conditional to inference prompt
After editing, the model stores the updated knowledge in the memory module $\bm{W}_{\text{m}}$, while the pre-trained knowledge is preserved in the original layer $\bm{W}_{0}$. To correctly respond to diverse prompts at inference stage, the model needs to retrieve the relevant knowledge from
@@
its memory. To do so, we propose the following two-step process (i) identifying the type of prompt during inference—whether it is an edited prompt, a rephrased version of an edited prompt, or a prompt irrelevant to the edits; and (ii) activating only the relevant knowledge in the memory. Specifically, it should retrieve the relevant knowledge distributed in corresponding combination of columns in $\bm{W}_{\text{m}}$ for the edited prompts and their rephrased variants, while deactivating the entire residual memory for irrelevant prompts for preserving locality.
During editing, we construct a database of sparse masks of each edited sample. Let $\bm{x}\in\mathcal{D}_{\textrm{edit}}$ be an edited prompt. Given that the layers prior to the introduced memory module are frozen, the application of the TopHash mechanism will yield the same sparse mask $\mathcal{M}(\bm{a}(\bm{x}))$ during both editing and inference, allowing the reliable identification of the prompt as part of the edit set by finding the exact match in the above database. Since transformers cluster semantically similar input prompts together due to their ability to generalize across paraphrases sentencebert , we expect that $\bm{x}$ should have activations, as well as TopHash masks, with high similarity w.r.t. its semantically rephrased variants. Conversely, semantically unrelated prompts will exhibit low similarity in both activation and mask space to the edited prompt $\bm{x}$.
@@
$\bm{x}_{\textrm{match}}$ $=\operatorname*{arg\,min}_{\bm{x}^{\prime}\in\mathcal{D}_{\textrm{edit}}}d_{\textrm{H}}\bigl{(}\mathcal{M}(\bm{a}(\bm{x})),\,\mathcal{M}(\bm{a}(\bm{x}^{\prime}))\bigr{)}$
We define the optimal overlapping ratio [^1] Note that the overlapping ratio is equivalent to the Hamming distance for two binary vectors with the exact same number of zeros and ones. between a prompt $\bm{x}$ and its closest edit $\bm{x}_{\textrm{match}}$ as:
@@
We intuitively classify the prompt type of the input $\bm{x}$ based on $R_{\textrm{match}}$ and a threshold $\tau$ and activate only the relevant knowledge from memory accordingly. Formally, the edited feedforward output at inference is then computed as:
$\text{FFN}_{\text{edited}}(\bm{a}(\bm{x}))=\begin{cases}\bm{W}_{0}\ \bm{a}(\bm{x})+\bm{W}_{\text{m}}\left(\mathcal{M}(\bm{a}(\bm{x}_{\textrm{match}}))\odot\bm{a}(\bm{x})\right)&\text{if }R_{\textrm{match}}(\bm{x})\in[\tau,1.0],\\ \bm{W}_{0}\ \bm{a}(\bm{x})&\text{if }R_{\textrm{match}}(\bm{x})\in[0,\tau)\end{cases}$ (4)
Specifically, if $R_{\textrm{match}}(\bm{x})\geq\tau$, we identify $\bm{x}$ as semantically relevant to an edited prompt $\bm{x}_{\textrm{match}}$, and apply the mask $\mathcal{M}(\bm{a}(\bm{x}_{\textrm{match}}))$ to its activations to activate the columns in $\bm{W}_{\text{m}}$ storing the relevant knowledge. This mechanism ensures that semantically rephrased prompts are aligned with their original edited counterparts in representation space, enabling precise retrieval of the relevant residual memory. As a result, it significantly improves generalization to paraphrased queries. Meanwhile, the residual memory is deactivated for prompts considered irrelevant to all edited prompts, i.e., $R_{\textrm{match}}(\bm{x})<\tau$, effectively preserving locality by relying solely on the pre-trained knowledge.
Prior approaches often require access to large corpora of irrelevant samples to enhance locality; either by sampling a different irrelevant sample during each edit (wang2024wise, ) or by performing a forward pass on a corpora of up to 100,000100{,}000 samples to regularize editing (memit, ; fang2025alphaedit, ). MEMOIR, on the other hand, eliminates the need for such large-scale irrelevant samples during editing and offers a more flexible alternative, as it relies on the introduced conditional knowledge activation mechanism in Equation 4 to decide whether to use the edited or pre-trained knowledge.